OpenStudy (anonymous):

If T=4st/s-t what is the value of s when t=5 and T=40?

OpenStudy (anonymous):

put the values in the equation ;) \[T = \left( \frac{ 4st }{ s-t } \right)\] t=5 , T=40 substituting the values, \[40 =\left( \frac{ 4s \left( 5 \right) }{ s - 5 } \right)\] \[40 = \frac{ 20s }{ s - 5 }\] \[40 \times \left( s - 5 \right) = 20s\] \[40s - 200 = 20s\] \[20s = 200\] \[s = \frac{ 200 }{ 20 }\] \[s = 10\]